# Do You Feel Comfortable? Detecting Hidden Conversational Escalation in AI Chatbots for Children

### v2

@@ 1 Introduction
Large Language Models (LLMs) are becoming deeply embedded in daily life as conversational agents, evolving beyond tools for information retrieval into companions for emotional support and social interaction vanhoffelen2025teens; hoffman2021parent. This trend is particularly pronounced among children and adolescents, a demographic that is prone to form parasocial relationships with AI chatbots xu2024growing; somerville2013teenage. Although beneficial, this dynamic introduces significant risks, as youth are uniquely vulnerable to manipulation due to neurodevelopmental changes mills2021inter; crone2012understanding; steinberg2005cognitive.
@@
Crucially, harmful conversational outcomes often emerge without explicitly toxic or abusive language. Even seemingly supportive or neutral responses can reinforce negative affect or normalize harmful states over repeated interactions, particularly for vulnerable userscheng2025social; sharma2023towards. This implicit harm remains largely invisible to safety mechanisms that rely on surface-level toxicity signals. Recent real-world incidents involving youth interactions with AI systems underscore the urgency of detecting these subtle yet consequential failures (see Appendix A).
Existing safety mechanisms typically focus on surface-level content moderation Yadav_Liu_Ortu_Ensafi_Jin_Mihalcea_2025 or single response toxicity classification deriu2021survey; LiZhang. Although recent frameworks have begun to address safety risks, they often rely on external classifiers or post-hoc analysis liu2025scales. These approaches may fail to capture the probabilistic momentum of a conversation, which is how a model’s response actively steers the user’s future emotional state wen2023unveiling.
@@ 2 Related Work
A common approach to conversational safety relies on external guardrail models that classify generated content into predefined risk categories, such as HateBERT caselli-etal-2021-hatebert, which adapts BERT-base to detect abusive and hateful language via supervised fine-tuning. Recent models like Llama-Guard are trained to identify policy violations and trigger refusals in conversational settings dubey2024llama3herdmodels. While effective for detecting explicit toxicity or rule-based violations, these models primarily operate as post-hoc binary classifiers over surface-level textual cues. As a result, they may struggle with implicit or context-dependent harms that lack overtly toxic markers.
@@
Our work aligns with white-box safety auditing. Methods like the LLM MicroscopeAzaria or first-token logit probingZhaoXuGupta demonstrate that models encode truthfulness and safety signals in their internal states razzhigaev2025llmmicroscopeuncoveringhiddenrole. GAUGE extends this by projecting these logits onto a learned affective space, allowing for interpretable monitoring of conversational drift without the need for external model or heavy retraining.
@@ 3.2 Stage 1: Risk Weight Calibration
We derive a reference weight vector $\lambda$ where positive values indicate a contribution to harm. We utilize the DiaSafety dataset, where dialogues are labeled as Safe (S=1S=-1) or Harmful ($S=+1$)sun2022safetyconversationalmodelstaxonomy. The complete pseudo-code for this calibration process is detailed in Algorithm 1 (see Appendix B).
@@ 4.1 Dataset
We use the DiaSafety testset sun2022safetyconversationalmodelstaxonomy, a dataset specifically curated to benchmark conversational safety in Human-AI interactions. DiaSafety consists of 10,000 dialogues where the risk is predominantly context-dependent.
@@
In addition, we evaluate the robustness of the attack on MinorBench khoo2025minorbench, a benchmark designed to assess whether safety mechanisms can be bypassed by benign linguistic but harmful instructions specially designed for children.
@@
We adopt Llama-3.1-8B-Instruct as the baseline model for all experiments dubey2024llama3herdmodels. All evaluations are conducted under standard autoregressive inference, using the model’s generated responses without additional rollouts or auxiliary sampling.
For affective grounding, we map the generated tokens to the eight primary categories of the NRC Emotion Lexicon. We benchmark our gradient-free framework against established external classifier baselines, including HateBERT caselli-etal-2021-hatebert, ToxicBERT dmonte2024generalizedoffensivelanguageidentification, and the Llama-Guard-3-8b dubey2024llama3herdmodels along with LLM termination baselines driven by safety system prompts (see Appendix C for the specific prompts used).
@@ 4.3 Results and Analysis
Table 2: Attack Success Rate (ASR) on MinorBench. An attack is considered successful if the model produces a safe response to an adversarial prompt. Lower is better.
@@
On MinorBench, GAUGE achieves a lower attack success rate compared to Llama-Guard-3-8B. While MinorBench prompts typically avoid explicit toxic or hateful expressions, many of them are harmful in a child safety context. In contrast, GAUGE operates by scoring responses along a continuous risk dimension, which allows it to flag harmful content even when surface-level indicators are absent. Even when GAUGE is instantiated as a binary classifier with a fixed threshold ($\tau=0.0$), it achieves an attack success rate of only 6% on MinorBench.
@@ 6 Conclusion
Appendix A Case Studies of Implicit Harm
@@
A.1 Character.AI: The “Sewell Setzer” Case
@@
A.2 Snapchat My AI: Contextual Blindness to Predation
@@
Appendix B Training Algorithm Pseudo-code
@@
Appendix C Experiment Prompts
C.1 Safety System Prompt (Refusal Baseline)
@@
Appendix D Discussion
GAUGE demonstrates that latent risk signals can be extracted directly from an LLM’s probability landscape.
Comparison with Guardrail Approaches: Unlike YouthSafe or VERA-MH which require training separate classifiers or defining complex clinical rubrics, GAUGE is unsupervised and model-intrinsic. It does not replace these systems but offers a complementary, lower-level signal (”affective velocity”) that can be used to trigger heavier, more expensive inspections.
False Positives and Therapeutic Empathy: A key challenge is distinguishing between harmful escalation and empathetic validation. A model reflecting a user’s sadness (e.g., ”I understand you are sad”) might trigger a high risk score in GAUGE if $\lambda$ heavily weights ”sadness.” This highlights the need for context-aware calibration. Future work should integrate pragmatic markers to separate therapeutic mirroring from maladaptive reinforcement.
Efficiency and Real-Time Feasibility: GAUGE requires only short rollouts ($T=5$), which incurs a small computational overhead compared to full chain-of-thought reasoning or external API calls. This makes it suitable for real-time monitoring in latency-sensitive applications like chat companions.

### v4

@@ 1 Introduction
Large Language Models (LLMs) are becoming deeply embedded in daily life as conversational agents, evolving beyond tools for information retrieval into companions for emotional support and social interaction Vanhoffelen et al. (2025); Hoffman et al. (2021). This trend is particularly pronounced among children and adolescents, a demographic that is prone to form parasocial relationships with AI chatbots Xu et al. (2024); Somerville (2013). Although beneficial, this dynamic introduces significant risks, as youth are uniquely vulnerable to manipulation due to neurodevelopmental changes Mills et al. (2021); Crone and Dahl (2012); Steinberg (2005).
@@
Crucially, harmful conversational outcomes often emerge without explicitly toxic or abusive language. Even seemingly supportive or neutral responses can reinforce negative affect or normalize harmful states over repeated interactions, particularly for vulnerable usersCheng et al. (2025); Sharma et al. (2023). This implicit harm remains largely invisible to safety mechanisms that rely on surface-level toxicity signals. Recent real-world incidents involving youth interactions with AI systems underscore the urgency of detecting these subtle yet consequential failures (see Appendix B).
Existing safety mechanisms typically focus on surface-level content moderation Yadav et al. (2025) or single response toxicity classification Deriu et al. (2021); Li et al. (2024). Although recent frameworks have begun to address safety risks, they often rely on external classifiers or post-hoc analysis Liu et al. (2025). These approaches may fail to capture the probabilistic momentum of a conversation, which is how a model’s response actively steers the user’s future emotional state Wen et al. (2023).
@@ 2 Related Work
A common approach to conversational safety relies on external guardrail models that classify generated content into predefined risk categories, such as HateBERT Caselli et al. (2021), which adapts BERT-base to detect abusive and hateful language via supervised fine-tuning. Recent models like Llama-Guard are trained to identify policy violations and trigger refusals in conversational settings Llama Team (2024). While effective for detecting explicit toxicity or rule-based violations, these models primarily operate as post-hoc binary classifiers over surface-level textual cues. As a result, they may struggle with implicit or context-dependent harms that lack overtly toxic markers.
@@
Our work aligns with white-box safety auditing. Methods like the LLM MicroscopeAzaria and Mitchell (2023) or first-token logit probingZhao et al. (2024) demonstrate that models encode truthfulness and safety signals in their internal states Razzhigaev et al. (2025). GAUGE extends this by projecting these logits onto a learned affective space, allowing for interpretable monitoring of conversational drift without the need for external model or heavy retraining.
@@ 3.2 Stage 1: Risk Weight Calibration
We derive a reference weight vector $\lambda$ where positive values indicate a contribution to harm. We utilize the DiaSafety dataset, where dialogues are labeled as Safe (S=1S=-1) or Harmful ($S=+1$)Sun et al. (2022). The complete pseudo-code for this calibration process is detailed in Algorithm 1 (see Appendix C).
@@ 4.1 Dataset
We use the DiaSafety testset Sun et al. (2022), a dataset specifically curated to benchmark conversational safety in Human-AI interactions. DiaSafety consists of 10,000 dialogues where the risk is predominantly context-dependent.
@@
In addition, we evaluate the robustness of the attack on MinorBench Khoo et al. (2025), a benchmark designed to assess whether safety mechanisms can be bypassed by benign linguistic but harmful instructions specially designed for children.
@@
We adopt Llama-3.1-8B-Instruct as the baseline model for all experiments Llama Team (2024). All evaluations are conducted under standard autoregressive inference, using the model’s generated responses without additional rollouts or auxiliary sampling.
We map the generated tokens of the NRC Emotion Lexicon. We benchmark our framework against established external classifier baselines, including HateBERT Caselli et al. (2021), ToxicBERT Dmonte et al. (2024), and the Llama-Guard-3-8b Llama Team (2024) along with LLM termination baselines driven by safety system prompts (see Appendix D for the specific prompts used).
@@ 4.3 Results and Analysis
Table 2: Attack Success Rate (ASR) on MinorBench. An attack is considered successful if the model fail to refuse to an adversarial prompt. Lower is better.
@@
Table 2 shows attack success rate of MinorBench, GAUGE achieves a lower attack success rate compared to Llama-Guard-3-8B. While MinorBench prompts typically avoid explicit toxic or hateful expressions, many of them are harmful in a child safety context. In contrast, GAUGE operates by scoring responses along a continuous risk dimension, which allows it to flag harmful content even when surface-level indicators are absent. Even when GAUGE is instantiated as a binary classifier with a fixed threshold ($\tau=0.0$), it achieves an attack success rate of only 6% on MinorBench.
@@ 6 Conclusion
7 Acknowledgement
This research is funded by the National Science Foundation under grant number 2125858. The authors would like to express their gratitude for the NSF’s support, which made this study possible.
Appendix A Statement on AI Assistance
AI-assisted tools were used to support the writing and editing of this manuscript. Specifically, large language models were employed to help improve clarity, organization, and grammatical correctness of the text. All scientific content, experimental design, implementation, analysis, and conclusions were developed and verified by the authors. The authors take full responsibility for the accuracy and integrity of the work.
Appendix B Case Studies of Implicit Harm
@@
B.1 Character.AI: The “Sewell Setzer” Case
@@
B.2 Snapchat My AI: Contextual Blindness to Predation
@@
Appendix C Training Algorithm Pseudo-code
@@
Appendix D Experiment Prompts
D.1 Safety System Prompt (Refusal Baseline)